# Interactive Distance Field Mapping and Planning to Enable Human-Robot Collaboration

### v1

@@ Abstract
Human-robot collaborative applications require scene representations that are kept up-to-date and facilitate safe motions in dynamic scenes. In this letter, we present an interactive distance field mapping and planning (IDMP) framework that handles dynamic objects and collision avoidance through an efficient representation. We define interactive mapping and planning as the process of creating and updating the representation of the scene online while simultaneously planning and adapting the robot’s actions based on that representation. Given depth sensor data, our framework builds a continuous field that allows to query the distance and gradient to the closest obstacle at any required position in 3D space. The key aspect of this work is an efficient Gaussian Process field that performs incremental updates and implicitly handles dynamic objects with a simple and elegant formulation based on a temporary latent model. In terms of mapping, IDMP is able to fuse point cloud data from single and multiple sensors, query the free space at any spatial resolution, and deal with moving objects without semantics. In terms of planning, IDMP allows seamless integration with gradient-based motion planners facilitating fast re-planning for collision-free navigation. The framework is evaluated on both real and synthetic datasets. A comparison with similar state-of-the-art frameworks shows superior performance when handling dynamic objects and comparable or better performance in the accuracy of the computed distance and gradient field. Finally, we show how the framework can be used for fast motion planning in the presence of moving objects. An accompanying video, code, and datasets are made publicly available [^3] https://uts-ri.github.io/IDMP.
@@ I Introduction
Human-robot collaboration (HRC) and other applications of robots in the field call for interactive representations to deal with dynamic and evolving scenes. For true collaboration in industrial settings, humans and robots physically share the same space, e.g. working jointly and simultaneously on the assembly of a product. The environment consisting of the human, robot, workspace, and objects such as tools and assembly components needs to be monitored by sensors and kept up-to-date in a timely manner. One approach is to create and continuously update an interactive representation of the changing scene which can be used for safe and effective robot motion planning without colliding with obstacles and human operators.
@@
Figure 1: Interactive generation of a distance and gradient field in an HRC setting. Top right: The depth image from an Intel Realsense camera. Bottom left: Coloured point cloud map of the scene and one horizontal slice of the field (red/blue means close/far to an object, arrows point away from the closest object).
@@
Recent representations that aim to fulfil some of the above-mentioned requirements have been proposed in the robotics literature [1, 2, 3]. Euclidean Distance Field (EDF), for instance, has recently become a promising representation suitable for direct collision-checking. As an example, the methods in [4, 5, 3] compute the projective or non-projective Truncated Signed Distance Field (TSDF) first, then propagate from TSDF to the free space for updating the EDF. Han et al. [1] propose to propagate from an occupancy map instead of the TSDF to maintain efficiency while preserving accuracy for the EDF computation. EDF representations also have the ability to dynamically update the changes via the so-called free space carving method. This method performs expensive ray-casting and progressive integration of the TSDF to update the map, resulting in free space that is only gradually cleared when objects move in the scene.
@@
This paper presents an interactive distance field mapping and planning (IDMP) framework aimed at dynamic scenes common in human-robot collaboration scenarios (see Fig. 1 as an example). We propose to build and maintain an up-to-date distance and gradient field using a Gaussian-Process-based method capable of integrating depth sensor measurements following our previous work in [6, 7]. Here, we present a simple and elegant method for dynamic updates and fusion of our Gaussian Process (GP) distance field by querying distances and gradients from a temporary latent GP distance field generated with only the points of the current frame. We name this latent representation the “Frustum Field”. The Frustum Field is able to deal with the discrepancy in the distance between surface points that move in the current frame without the need for ray-tracing. It also allows seamless fusion with previously mapped distance and gradient field.
@@
To overcome the computational complexity of GP-based approaches, our framework maintains the fusion and updates in an Octree structure that runs online. Furthermore, our method implicitly handles dynamic objects via the information provided by the Frustum Field. The fused representation is continuous allowing querying of the Euclidean distance to the nearest surface and its gradient at an arbitrary spatial resolution. This property enables seamless integration with a gradient-based motion planning framework. We experimentally evaluate IDMP’s performance in both static and dynamic scenes and benchmark it against state-of-the-art algorithms.
@@
  • An interactive distance field mapping and planning (IDMP) framework that generates a continuous Euclidean distance and gradient field online, and is integrated with a gradient-based planner for 3D collision-free navigation.
  • An efficient open-sourced ROS-based implementation of the full IDMP framework.  [^1] Code will be made publicly available at https://github.com/UTS-RI/IDMP upon paper acceptance.
@@ II Related Work
EDFs are useful representations for a wide range of tasks which require distance-to-obstacle information at different points in space. For example, motion planners such as gradient-based trajectory optimisers [8, 9] utilise EDFs and their gradients to perturb trajectories smoothly in order to avoid obstacles. EDFs can be computed by solving the Eikonal equation. A major challenge in exploiting the Eikonal equation to recover the ESDF is that it is non-linear and hyperbolic. Therefore, it is difficult to solve directly, and specialised numerical approaches are often required, including discrete methods [10] such as fast marching or label-correcting to propagate the distance field through a grid.
Inspired by [10], Voxblox [4] proposes an approach to compute the ESDF based on the projective TSDF through a wavefront propagation algorithm. Voxfield [5] and VDBblox [3] adopt the use of a non-projective TSDF to improve the accuracy. FIESTA [1] proposes to use an occupancy map instead of TSDF and a novel data structure to maintain efficiency. VED-EDT uses a distance transform function and VDB data structure [11, 12] to represent the EDF hierarchically [13]. These EDF representations both have the capability to dynamically update the changes via the so-called free space carving method, which updates every voxel along the grouped rays. This method requires performing expensive ray-casting and progressive integration and weighting of the TSDF to update the map, resulting in free space that is only gradually cleared when objects move in the scene. There are also frameworks that explicitly deal with dynamic objects [14]. However, the dynamic object handling is decoupled from mapping. Our framework deals implicitly with dynamic objects and their surfaces are considered for collision avoidance, thus tightly coupled with the estimated representation.
Over the last few years, there have been numerous works exploring the potential of learning approaches such as neural networks for implicit signed distance representations. In particular, inspired by [15], the authors in [16] propose to use an implicit neural representation to constrain the learning through the Eikonal regularisation, which approximates the SDF for points far away from the surfaces. Recently proposed in [17], the authors learn an approximation of EDF from Neural Radiance Fields (NeRFs) with occupancy inference. Similarly, inspired from [16], iSDF [18] is proposed to use a neural signed distance field for mapping. To enable the capability for dynamic scenes, D-NeRF considers time as an additional input to the system and splits the learning process into two main stages: one that encodes the scene into a canonical space and another that maps this canonical representation into the deformed scene at a particular time. While these approaches have benefits such as a continuous representation and lower memory cost, they are expensive to compute and require powerful high-end hardware with a GPU.
@@ III-A Gaussian Process Distance Field
$f\sim\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}^{\prime})),$ (1)
where $k(\mathbf{x},\mathbf{x}^{\prime})$ is the covariance function that controls the properties of $f$. Suppose we have a set of testing points $\mathbf{X_{*}}=\{\mathbf{x_{*}}_{q}\}_{q=1}^{Q}\in\mathbb{R}^{D}$. At the testing locations, we can express the joint distribution of the function values and the observed target values using the kernel function,
$\left[\begin{array}[]{l}\mathbf{y}\\ \mathbf{f}_{*}\end{array}\right]=\mathcal{N}\left(\mathbf{0},\left[\begin{array}[]{ll}\mathbf{K}(\mathbf{X},\mathbf{X})+\sigma_{f}^{2}\mathbf{I}&\mathbf{K}\left(\mathbf{X},\mathbf{X_{*}}\right)\\ \mathbf{K}\left(\mathbf{X_{*}},\mathbf{X}\right)&\mathbf{K}\left(\mathbf{X_{*}},\mathbf{X_{*}}\right)\end{array}\right]\right).$ (2)
$\mathbf{f}_{*}$ is the vector of predicted function values at $\mathbf{X_{*}}$, $\mathbf{I}$ is the identity matrix, $\sigma_{f}^{2}$ is the variance of observation noise and $\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)$ is the $J\times J$ covariance matrix between input points. Thus, by conditioning Eq. (2), the posterior distribution of $f$ at an arbitrary testing point $\mathbf{x}_{*}$ is given by $f(\mathbf{x}_{*})\sim\mathcal{N}(\hat{f}\left(\mathbf{x_{*}}\right),\mathbb{V}\hat{f}\left(\mathbf{x_{*}}\right))$, where the predictive mean and variance are:
@@
$\mathbf{k}(\mathbf{x_{*}},\mathbf{X})$ is the vector of covariances between the $J$ input points and the testing point. $k\left(\mathbf{x}_{*},\mathbf{x}_{*}\right)$ is the covariance function value of the testing point.
Let us now briefly introduce the so-called reverting GP distance field originally presented in [7]. Consider a surface $\mathcal{S}$ in a Euclidean space $\mathbb{R}^{D}$, and $\mathbf{X}=\{\mathbf{x}_{j}\}_{j=1}^{J}\in\mathbb{R}^{D}$ a set of discrete observations of $S$ as training points. By modelling the occupancy $o(\mathbf{x}):\mathbb{R}^{D}\mapsto\mathbb{R}$ of the space with a GP as $o\sim\mathcal{GP}\left(0,k(\mathbf{x},\mathbf{x}^{\prime})\right)$, it is possible to infer the surface occupancy $\hat{o}(\mathbf{x}_{*})$ at any location in the space using Eq. (3). In other words, the covariance kernel ${{k}\left(\mathbf{x},\mathbf{x}^{\prime}\right)}$ is a monotonic function of the distance $\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert$ between $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
@@
The distance field ${\hat{d}}(\mathbf{x}_{*})$ given any location $\mathbf{x}_{*}$ in the space is obtained by applying a reverting function $r$ to the occupancy field as
@@
with $\text{with }\ \tilde{{{k}}}(\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert)$ $={{k}\left(\mathbf{x},\mathbf{x}^{\prime}\right)}.$
Considering the square exponential kernel ${{k}\left(\mathbf{x},\mathbf{x}^{\prime}\right)}=\sigma^{2}\exp\left(-\frac{\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert{}^{2}}{2l^{2}}\right)$, substituting the reverting function of square exponential kernel into Eq. 6 gives us:
@@
III-B Linear Operator for Gradient Field
The derivative of a GP is another GP. One common solution to infer the gradient field is to have surface normals as input for GP modelling. However, this requires inverting a computational-heavy joint covariance function of $k(\mathbf{x},\mathbf{x}^{\prime})$ with the partial derivatives of $k(\mathbf{x},\mathbf{x}^{\prime})$ at $\mathbf{x}$ and $\mathbf{x}^{\prime}$ [33]. Instead, we propose to use the linear operator to infer the gradient along with the accurate distance field without the normal as input. Applying the linear operator to Eq. 5, we have
@@ III-B Linear Operator for Gradient Field
where $\nabla\mathbf{k}(\mathbf{x_{*}}\mathbf{X})$ is the partial derivative of $\mathbf{x}$ (gradient matrix) and $\nabla k_{\mathbf{x}_{*}\mathbf{x}_{*}}\nabla$ is the second-order partial derivatives of $\mathbf{x}$ and $\mathbf{x}^{\prime}$ (Hessian matrix). Taking the gradient of both sides of Eq. 8 with respect to the distance, the gradient of ${\hat{d}}(\mathbf{x}_{*})$ is in the same direction as the gradient of $\hat{o}(\mathbf{x}_{*})$, subject to a scaling factor, so that $\nabla{\hat{d}}(\mathbf{x}_{*})\approx\nabla\hat{o}(\mathbf{x}_{*})$.
Figure 2: System diagram of IDMP. We first model the temporary latent Frustum Field (blue) using only $\mathcal{P}_{\{i\}}$ as training points. All prior training points $\mathcal{P}_{\{0,...,i-1\}}$ in the Fused Field (yellow) are then passed to the Frustum Field. Given the sensor pose, the Frustum Field selects from $\mathcal{P}_{\{0,...,i-1\}}$ the points that are within the frustum area $\mathcal{P}_{f\{0,...,i-1\}}$ and returns the inferred values of $\hat{d}_{f}$, $\nabla\hat{d}_{f}$ to the Fused Field. These distances and gradients are used to perform fusion and dynamic update by updating the training points that model the Fused Field. The path planner then queries for $\hat{d}_{o}$, $\nabla\hat{d}_{o}$ in order to compute and adapt its motion plans in response to a changing map.

### v3

@@ Abstract
Human-robot collaborative applications require scene representations that are kept up-to-date and facilitate safe motions in dynamic scenes. In this letter, we present an interactive distance field mapping and planning (IDMP) framework that handles dynamic objects and collision avoidance through an efficient representation. We define interactive mapping and planning as the process of creating and updating the representation of the scene online while simultaneously planning and adapting the robot’s actions based on that representation. The key aspect of this work is an efficient Gaussian Process field that performs incremental updates and handles dynamic objects reliably by identifying moving points via a simple and elegant formulation based on queries from a temporary latent model. In terms of mapping, IDMP is able to fuse point cloud data from single and multiple sensors, query the free space at any spatial resolution, and deal with moving objects without semantics. In terms of planning, IDMP allows seamless integration with gradient-based reactive planners facilitating dynamic obstacle avoidance for safe human-robot interactions. Our mapping performance is evaluated on both real and synthetic datasets. A comparison with similar state-of-the-art frameworks shows superior performance when handling dynamic objects and comparable or better performance in the accuracy of the computed distance and gradient field. Finally, we show how the framework can be used for fast motion planning in the presence of moving objects both in simulated and real-world scenes. An accompanying video, code, and datasets are made publicly available [^3] https://uts-ri.github.io/IDMP.
@@ I Introduction
Human-robot collaboration (HRC) and other applications of robots in the field call for interactive representations to deal with dynamic and evolving scenes. For true collaboration in industrial settings, humans and robots physically share the same space, e.g. working jointly and simultaneously on the assembly of a product. The environment consisting of the human, robot, workspace, and objects such as tools and assembly components needs to be monitored by sensors and kept up-to-date in a timely manner. One approach is to create and continuously update an interactive representation of the changing scene, which is useful for safe and effective robot motion planning around human operators.
@@
Figure 1: Interactive generation of a distance and gradient field in an HRC setting. Top right: Depth image from an Intel Realsense camera. Bottom left: Coloured point cloud and a horizontal slice of the field (red/blue means close/far to the nearest object with arrows pointing away from it).
@@
Recent representations that aim to fulfil some of the above-mentioned requirements have been proposed in the robotics literature [1, 2]. Euclidean Distance Field (EDF), for instance, has recently become a promising representation suitable for direct collision-checking. As an example, the methods in [3, 4, 2] compute the projective or non-projective Truncated Signed Distance Field (TSDF) first, then propagate from TSDF to the free space for updating the EDF. Han et al. [1] propose to propagate from an occupancy map instead of the TSDF to maintain efficiency while preserving accuracy for the EDF computation. EDF representations also have the ability to dynamically update the changes via the so-called free space carving method. This method performs expensive ray-casting and progressive integration of the TSDF to update the map, resulting in free space that is only gradually cleared when objects move in the scene.
@@
This paper presents an interactive distance field mapping and planning (IDMP) framework aimed at dynamic scenes common in human-robot collaboration scenarios (see Fig. 1 as an example). By extending our previous work in [5, 6], we focus to build and maintain an efficient and up-to-date distance and gradient field using a Gaussian-Process-based method capable of integrating depth sensor measurements. Here, we present a simple and elegant method catering both dynamic updates and seamless fusion by querying distances and gradients from a temporary latent GP distance field generated with only the points of the current frame. We name this latent representation the “Frustum Field”. The Frustum Field is able to deal with the discrepancy in the distance between surface points that move in the current frame without the need for ray tracing. It also allows fusion with previously mapped distance and gradient fields.
@@
To overcome the computational complexity of GP-based approaches, our framework maintains the updates and fusion in an Octree structure that runs online. Furthermore, our method handles dynamic objects reliably via the information provided by the Frustum Field. The fused representation is continuous, allowing querying of the Euclidean distance to the nearest surface and its gradient at an arbitrary spatial resolution. To demonstrate the potential of our efficient and differentiable fused distance and gradient fields for human-robot interaction tasks, we integrate IDMP with a gradient-based motion planner as well as a local reactive behaviour. We experimentally evaluate IDMP’s performance in both static and dynamic scenes and benchmark it against state-of-the-art algorithms and demonstrate its capabilities in real-world experiments for a robot arm interacting with a human.
@@
  • An interactive distance field mapping and planning (IDMP) framework that generates a continuous Euclidean distance and gradient field online, and is integrated with gradient-based reactive motion and path planner for 3D collision-free interactions.
  • An efficient open-sourced ROS-based implementation of the full IDMP framework.  [^1] Code is made publicly available at https://github.com/UTS-RI/IDMP.
@@ II Related Work
EDFs are useful representations for a wide range of tasks which require distance-to-obstacle information at different points in space. For example, motion planners such as gradient-based trajectory optimisers [7, 8] utilise EDFs and their gradients to perturb trajectories smoothly in order to avoid obstacles. EDFs can be computed by solving the Eikonal equation. A major challenge in exploiting the Eikonal equation to recover the ESDF is that it is non-linear and hyperbolic. Therefore, it is difficult to solve directly, and specialised numerical approaches are often required, including discrete methods such as fast marching or label-correcting to propagate the distance field through a grid.
Voxblox [3] proposes an approach to compute the ESDF based on the projective TSDF through a wavefront propagation algorithm. Voxfield [4] and VDBblox [2] adopt the use of a non-projective TSDF to improve the accuracy. FIESTA [1] proposes to use an occupancy map instead of TSDF and a novel data structure to maintain efficiency. VED-EDT uses a distance transform function to represent the EDF hierarchically [9]. These EDF representations both have the capability to dynamically update the changes via the so-called free space carving method, which updates every voxel along the grouped rays. This method requires performing expensive ray-casting and progressive integration and weighting of the TSDF to update the map, resulting in free space that is only gradually cleared when objects move in the scene. There are also frameworks that explicitly deal with dynamic objects [10]. However, the dynamic object handling is decoupled from mapping. Our framework deals reliably with dynamic objects and their surfaces are considered for collision avoidance, thus tightly coupled with the estimated representation.
Over the last few years, there have been numerous works exploring the potential of learning approaches such as neural networks for implicit signed distance representations. In particular, inspired by [11], the authors in [12] propose to use an implicit neural representation to constrain the learning to approximate the SDF. Recently proposed in [13], the authors learn an approximation of EDF from Neural Radiance Fields (NeRFs) with occupancy inference. Similarly, inspired from [12], iSDF [14] is proposed to use a neural signed distance field for mapping. For dynamic scenes, D-NeRF considers time as an additional input to the system and splits the learning process into two main stages: one that encodes the scene into a canonical space and another that maps this canonical representation into the deformed scene at a particular time. While these approaches have benefits such as a continuous representation and lower memory cost, they are expensive to compute and require powerful GPUs.
@@ III-A Gaussian Process Distance Field
$f\sim\mathcal{GP}(0,k_{\mathbf{x}\mathbf{x}^{\prime}}),$ (1)
where $k_{\mathbf{x}\mathbf{x}^{\prime}}$ is the covariance function that controls the properties of $f$. Suppose we have a set of testing points $\mathbf{X_{*}}=\{\mathbf{x_{*}}_{q}\}_{q=1}^{Q}\in\mathbb{R}^{D}$. At the testing locations, we can express the joint distribution of the function values and the observed target values using the kernel function,
$\left[\begin{array}[]{l}\mathbf{y}\\ \mathbf{f}_{*}\end{array}\right]=\mathcal{N}\left(\mathbf{0},\left[\begin{array}[]{ll}\mathbf{K}_{\mathbf{X}\mathbf{X}}+\sigma_{f}^{2}\mathbf{I}&\mathbf{K}_{\mathbf{X}\mathbf{X_{*}}}\\ \mathbf{K}_{\mathbf{X_{*}}\mathbf{X}}&\mathbf{K}_{\mathbf{X_{*}}\mathbf{X_{*}}}\end{array}\right]\right).$ (2)
$\mathbf{f}_{*}$ is the vector of predicted function values at $\mathbf{X_{*}}$, $\mathbf{I}$ is the identity matrix, $\sigma_{f}^{2}$ is the variance of observation noise and $\mathbf{K}_{\mathbf{X}\mathbf{X}}$ is the $J\times J$ covariance matrix between input points. By conditioning Eq. (2), the posterior distribution of $f$ at an arbitrary testing point $\mathbf{x}_{*}$ is given by $f(\mathbf{x}_{*})\sim\mathcal{N}(\hat{f}\left(\mathbf{x_{*}}\right),\mathbb{V}\hat{f}\left(\mathbf{x_{*}}\right))$, where predictive mean and variance are:
@@
$\mathbf{k}_{\mathbf{x_{*}}\mathbf{X}}$ is the vector of covariances between the $J$ input points and the testing point. $k_{\mathbf{x}_{*}\mathbf{x}_{*}}$ is the covariance function value of the testing point.
Let us now briefly introduce the so-called reverting GP distance field originally presented in [6]. Consider a surface $\mathcal{S}$ in a Euclidean space $\mathbb{R}^{D}$, and $\mathbf{X}=\{\mathbf{x}_{j}\}_{j=1}^{J}\in\mathbb{R}^{D}$ a set of discrete observations of $S$ as training points. By modelling the occupancy $o(\mathbf{x}):\mathbb{R}^{D}\mapsto\mathbb{R}$ of the space with a GP as $o\sim\mathcal{GP}\left(0,k_{\mathbf{x}\mathbf{x}^{\prime}}\right)$, it is possible to infer the surface occupancy $\hat{o}(\mathbf{x}_{*})$ at any location in the space using Eq. (3). In other words, the covariance kernel $k_{\mathbf{x}\mathbf{x}^{\prime}}$ is a monotonic function of the distance $\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert$ between $\mathbf{x}$ and $\mathbf{x}^{\prime}$.
@@
The distance field ${\hat{d}}(\mathbf{x}_{*})$ given any location $\mathbf{x}_{*}$ is obtained by applying a reverting function $r$ to the occupancy field as
@@
with $\text{with }\ \tilde{{{k}}}(\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert)$ $=k_{\mathbf{x}\mathbf{x}^{\prime}}.$
Considering the square exponential kernel $k_{\mathbf{x}\mathbf{x}^{\prime}}=\sigma^{2}\exp\left(-\frac{\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert{}^{2}}{2l^{2}}\right)$, substituting the reverting function of it into Eq. 6 gives us:
@@
III-B Nabla Operator for Gradient Field
The derivative of a GP is a linear operation that produces another GP. One common solution to infer the gradient field is to have surface normals as input for GP modelling. However, this requires inverting a computational-heavy joint covariance function of $k_{\mathbf{x}\mathbf{x}^{\prime}}$ with the partial derivatives of $k_{\mathbf{x}\mathbf{x}^{\prime}}$ at $\mathbf{x}$ and $\mathbf{x}^{\prime}$ [30]. Instead, we propose to use the Nabla operator to infer the gradient along with the accurate distance field without the normal as input [31]. Applying this linear operator to Eq. 5, we get
@@ III-B Nabla Operator for Gradient Field
where $\nabla\mathbf{k}_{\mathbf{x_{*}}\mathbf{X}}$ is the partial derivative with respect to $\mathbf{x}$ [32]. Taking the gradient of both sides of Eq. 8 with respect to the distance shows that the gradient of ${\hat{d}}(\mathbf{x}_{*})$ points in the same direction as the gradient of $\hat{o}(\mathbf{x}_{*})$, subject to a scaling factor, so that $\nabla{\hat{d}}(\mathbf{x}_{*})\approx\nabla\hat{o}(\mathbf{x}_{*})$.
Figure 2: System diagram of IDMP. We first model the temporary latent Frustum Field (blue) using only $\mathcal{P}_{\{i\}}$ as training points. All prior training points $\mathcal{P}_{\{0,...,i-1\}}$ in the Fused Field (yellow) are then passed to the Frustum Field. Given the sensor pose, the Frustum Field selects from $\mathcal{P}_{\{0,...,i-1\}}$ the points that are within the frustum area $\mathcal{P}_{f\{0,...,i-1\}}$ and returns the inferred values of $\hat{d}_{f}$, $\nabla\hat{d}_{f}$ to the Fused Field. These distances and gradients are used to perform fusion and dynamic updates by updating the training points that model the Fused Field. The path planner then queries for $\hat{d}_{o,s}$, $\nabla\hat{d}_{o,s}$ in order to compute and adapt its motion plans in response to a changing map.